# captionUnsupported document class

### v1

@@ Abstract
Equipping autonomous robots with the ability to navigate safely and efficiently around humans is a crucial step toward achieving trusted robot autonomy. However, generating robot plans while ensuring safety in dynamic multi-agent environments remains a key challenge. Building upon recent work on leveraging deep generative models for robot planning in static environments, this paper proposes CoBL-Diffusion, a novel diffusion-based safe robot planner for dynamic environments. CoBL-Diffusion uses Control Barrier and Lyapunov functions to guide the denoising process of a diffusion model, iteratively refining the robot control sequence to satisfy the safety and stability constraints. We demonstrate the effectiveness of the proposed model using two settings: a synthetic single-agent environment and a real-world pedestrian dataset. Our results show that CoBL-Diffusion generates smooth trajectories that enable the robot to reach goal locations while maintaining a low collision rate with dynamic obstacles.
@@ I Introduction
Diffusion models [14, 15] have emerged as a powerful class of deep generative models, achieving remarkable success in generating high-quality images and synthesizing speech [16, 17], and more recently, in robot trajectory generation [18, 19, 20, 21]. However, current diffusion-based robot trajectory generation methods still lack explicit safety assurances in dynamic environments, prohibiting their application in safety-critical scenarios, such as those involving human interactions. In this work, we investigate the use of diffusion models as the foundation for robot navigation planning in dynamic multi-agent environments and present a method that integrates control-theoretic safety and stability constraints into the diffusion model framework.
@@
    We propose CoBL-Diffusion, a novel diffusion-based robot planner for dynamic environments that incorporates control-theoretic safety and stability constraints. The diffusion process is guided by the gradient of reward functions derived from the CBF and CLF.
@@
    Our method ensures dynamic consistency between the control sequence and the resulting states by explicitly integrating generated controls through system dynamics to obtain the states. Maintaining dynamic consistency during the diffusion steps is crucial, as the CBF and CLF reward computations in CoBL-Diffusion require the state trajectory to be consistent with the control inputs.
@@ II Related work
Diffusion Models for Planning. Diffusion models, introduced by [14] and further developed by [15], are a class of deep generative models that represent data generation as an iterative denoising process. We provide a brief overview of diffusion models in Section III-A. The Diffuser from [18] represents an innovative application of diffusion models to robot trajectory generation. The Diffuser combines the planning and sampling processes into a unified framework via iterative denoising of trajectories, thus enabling the generation of plans that are adaptable to various reward structures and constraints. There are follow-up works that aim to influence the diffusion process to produce trajectories that satisfy some desired properties. Work similar in spirit to our proposed method, SafeDiffuser [32] extends the application of diffusion models to safety-critical tasks by incorporating CBFs into the denoising diffusion procedure to ensure the generated trajectories satisfy safety constraints. This CBF guarantees forward invariance of the diffusion procedure rather than the conventional forward invariance in a time domain addressed in this work. Another recent work [19] proposes guided conditional diffusion in the context of simulating human driving behaviors in driving simulators. They use signal temporal logic specifications to encode desired behaviors such as goal-reaching and staying within speed limits.
@@ III-A Diffusion Models
where $\boldsymbol{\epsilon}$ is Gaussian noise, $\boldsymbol{\epsilon}_{\theta}$ is the neural network predicting the noise, and $\bar{\alpha}_{i}=\prod_{j=1}^{i}(1-\beta_{j})$ is the cumulative product of the noise schedule. To generate new samples, one starts from noise and iteratively applies (2) until for some predetermined number of iterations.
@@ III-C Control Barrier Function (CBF)
$\sup_{\mathbf{u}\in\mathcal{U}}\mathopen{}\mathclose{{}\left[L_{f}h(\mathbf{x})+L_{g}h(\mathbf{x})\mathbf{u}+\alpha(h(\mathbf{x}))}\right]\geq 0,\ \forall\mathbf{x}\in\mathcal{D}.$ (9)
@@ IV-B1 U-Net Conditioning on Trajectories
To maintain dynamic consistency, our model generates only control inputs, that is, compared to (4), we have $\boldsymbol{\tau}=\mathbf{u}=[\mathbf{u}_{1},\ \ldots,\ \mathbf{u}_{T}]$, and the states are obtained by feeding the controls into the system dynamics (13) (see the blue box in Fig. 2 left), rather than simultaneously predicting control inputs and states. While Diffuser conditions the goal by replacing the terminal state, the same inpainting method cannot be applied to CoBL-Diffusion since it predicts controls only. Therefore, the states to be satisfied are fed into each resolution of the U-Net [41] (see orange boxes in Fig. 2) to condition the generated controls for the next denoising step. During the training of the proposed model, ground truth states are corrupted with the same noise as the control inputs in the forward diffusion process (1) before feeding them into the model. This is because, during the generation process, the states obtained by integrating noisy controls with the system dynamics (13) are noisy.
@@ IV-B3 Temporal Weighting for Covariance
The behavior of a generated trajectory depends on the contrl sequence by integrating the dynamics. As a result, errors in control inputs accumulate over time, meaning errors at earlier timesteps can lead to large errors in states later in the horizon. Therefore, denoising the later control sequence has little effect if the earlier control inputs have not converged. To encourage the convergence of the earlier control sequence, the covariances of the forward diffusion (1) and reverse diffusion (2) are weighted by $\mathbf{V}$ as follows:
@@
where $v_{0},\ldots,v_{T}$ are scheduled to increase monotonically.
@@ IV-C Reward Function
As introduced in Section III-B, a user-defined reward function can be used to guide the reverse diffusion process (5) to generate outputs with some desirable behavior. In this work, the reverse diffusion process can be guided by multiple reward functions simultaneously, therefore the optimality of the trajectory is denoted using the set of $K$ reward functions $W_{k}(\mathbf{x}_{t},\mathbf{u}_{t})$ as $p(O_{t}=1)=\textrm{exp}\mathopen{}\mathclose{{}\left(\sum_{k=1}^{K}W_{k}(\mathbf{x}_{t},\mathbf{u}_{t})}\right)$. Then, we modify the reverse diffusion process as follows:
@@ IV-C1 Control Barrier Function Reward
Therefore, if the reward function $W_{\mathrm{cbf}}$ can be made positive at the end of the denoising process, it is guaranteed that the generated control sequence will avoid a collision with the obstacle. The gradient of this reward function with respect to the control input $\mathbf{u}$ is computed as follows:
@@ IV-C2 Control Lyapunov Function Reward
The reward function based on Control Lyapunov Functions (CLFs) in Section III-D guides the reverse diffusion process to generate a control sequence that ensures convergence to a given goal $\mathbf{x}_{g}$. Similar to CBF reward, we assume that the dynamics of the robot is (7).
Suppose $h_{\mathrm{clf}}$ is a CLF, then, from Theorem 2, control inputs $\mathbf{u}$ that make the following reward function $W_{\mathrm{clf}}$ positive will drive the state of the robot $\mathbf{x}$ to $\mathbf{x}_{g}$:
@@
Therefore, guiding the reverse diffusion process to make this reward function $W_{\mathrm{clf}}$ positive ensures convergence to the goal. The gradient of this reward function with respect to the control input $\mathbf{u}$ is computed as follows:
@@ IV-D Conditional Motion Planning with Reverse Diffusion
First, we observe start $\mathbf{x}_{s}$ and goal $\mathbf{x}_{g}$ of the trajectory, then initialize the trajectory $\mathbf{x}^{N}$, where $N$ is the number of diffusion steps. If there is prior information about the environment, we can initialize based on that; otherwise, we simply initialize the trajectory as a straight line connecting the start and goal points. Next, we sample $\mathbf{u}^{N-1}$ conditioned on $\mathbf{x}^{N}$ according to the following distribution:
@@
Then, the control inputs are fed into the system dynamics (13) to obtain the trajectory $\mathbf{x}^{N-1}$. At this point, it is important to emphasize that, unlike conventional diffusion model-based trajectory generation, the control sequence and states are consistent. Then, the trajectory is evaluated by the reward functions defined in Section IV-C. If each reward function is positive at a certain time step, it indicates the trajectory already satisfies each condition at that time. Therefore, the gradient of the reward function is masked with 00 at that point. Then, the guided controls $\hat{\mathbf{u}}^{i}$ are obtained by adding the gradient to the denoised controls $\mathbf{u}^{i}$ and the updated states $\hat{\mathbf{x}}^{i}$ are obtained using $\hat{\mathbf{u}}^{i}$. Before fed the updated states, initial and terminal states are replaced with the given start $\mathbf{x}_{s}$ and goal $\mathbf{x}_{g}$ to condition the controls. Then, the controls $\hat{\mathbf{u}}^{i}$ are denoised based on the states $\hat{\mathbf{x}}^{i}$. This approach aims to iteratively generate controls that gradually satisfy the conditions, instead of employing a common approach of solving a quadratic program (QP) to enforce satisfaction of CBF/CLF constraints which can be computationally expensive over many diffusion and timesteps. Since there is no need to solve a QP, there is no concern about the feasibility of the trajectories satisfying the constraints. When using multiple reward functions, the prioritization of conditions can be managed by adjusting their coefficients. For example, safety constraints should be emphasized over other constraints, thus necessitating a higher coefficient.
@@ V Experiments and Discussion
In this section, we validate the effectiveness of the proposed CoBL-Diffusion to synthesize safe and goal-reaching controllers in dynamic environments. First, we investigate the characteristics of controllers resulting from different reward functions in a single-agent environment. Then, we validate the efficacy in a multi-agent environment using the UCY pedestrian dataset [45].
@@ V-A Model Training and Environment Setup
We trained CoBL-Diffusion using the ETH pedestrian dataset [46] for 500 epochs on 276,874 8-second trajectories. The architecture and hyperparameters of the diffusion model are based on an open-source implementation [^3] https://github.com/jannerm/diffuser. The covariance weight (19) was set to increase linearly from $0.5$ to $1$. We evaluated the proposed model in two simulation environments created using Trajdata [47]. The first simulation environment is a toy environment with a single moving obstacle. The second environment is from the UCY pedestrian dataset [45] for challenging multi-agent yet feasible settings.
We assumed the robot follows single integrator dynamics, $\dot{\mathbf{x}}=\mathbf{u}$, noting that our algorithm can easily be extended to other types of dynamics. We use the following CBF and CLF:
@@
where $\mathbf{x}=[x,y]$, $\mathbf{x}_{q}=[x_{q},y_{q}]$ represent the states of the robot and the $q$-th obstacle respectively, $r$ is a barrier radius, and $\mathbf{x}_{g}=[x_{g},y_{g}]$ is a given goal point. Given the dynamics and CBF/CLF, we have a relative degree one system.
@@ V-B Comparison Methods and Evaluation Metrics
We compare CoBL-Diffusion (CoBL) with the baseline method CBF-QP, which solves a QP with CBF at every timestep given a nominal straight line trajectory to the goal, and Velocity Obstacle (VO) [48], which is a geometric approach for collision avoidance that uses the relative velocity between a robot and an obstacle to identify potential collision velocities. Variants of the proposed CoBL-Diffusion are also investigated: dCoBL-Diffusion (dCoBL), which uses discrete-time CBF and CLF rewards; CoB-Diffusion (CoB), which only uses a CBF reward;
DCoL-Diffusion (DCoL), which uses CLF and a distance reward $W_{\mathrm{dist}}$ instead of a CBF reward:
@@
where $r$ is a barrier radius. Comparing the CBF reward, the distance reward considers only the distance to obstacles and ignores the dynamics. Here, the coefficients for (discrete-time) CBF and CLF rewards are fixed at $0.3$ and $0.1$, and for distance reward is fixed at $0.3$. In this experiment, we set the barrier radius to $1\ m$ and the collision radius to $1\ m$. For a clear comparison, the reward functions for collision avoidance consider only the nearest obstacle, although considering multiple obstacles could potentially improve performance and is deferred for future work.
Each method is compared against three planning metrics: collision rate, goal-reaching, and smoothness. The collision rate is calculated as the ratio of simulations that violate the collision radius among all simulations. The goal-reaching is evaluated by the root squared error between the goal and the final point if the plan is followed. The smoothness is measured by the worst-case root squared difference in control inputs (in these experiments, velocity).
@@ V-C Safe Planning in Single-Agent Environment
CoBL $W_{\mathrm{CBF}}$ $W_{\mathrm{CLF}}$ $\mathbf{0}$ $\mathbf{0.18\pm 0.02}$ $\mathbf{0.03\pm 0.01}$ $\mathbf{0.5}$ $\mathbf{0.41\pm 0.20}$ $\mathbf{0.07\pm 0.03}$
CoB $W_{\mathrm{CBF}}$ 00 % $0.20\pm 0.10$ $0.18\pm 0.02$ $20.0$ $3.60\pm 1.74$ $0.20\pm 0.05$
dCoBL $W_{\mathrm{dCBF}}$ $W_{\mathrm{dCLF}}$ $9$ $0.16\pm 0.08$ $0.37\pm 0.01$ $48.0$ $0.38\pm 0.16$ $0.05\pm 0.01$
DCoL $W_{\mathrm{dist}}$ $W_{\mathrm{CLF}}$ $8$ $0.12\pm 0.08$ $0.24\pm 0.02$ $42.5$ $1.10\pm 1.89$ $0.15\pm 0.19$
CoBL- $W_{\mathrm{CBF}}$ $W_{\mathrm{CLF}}$ $8.5$ $0.48\pm 0.34$ $0.08\pm 0.04$
CBF-QP 00 % $0.13$ $1.08$ $62.5$ $0.07\pm 0.02$ $0.50\pm 0.42$
@@
(b) CBF-QP.
(c) VO.
(d) dCoBL-Diffusion.
(e) CoB-Diffusion.
(f) DCoL-Diffusion.
Figure 3: Generated trajectories in the single-agent environment with various planning methods. Red trajectories experience collision.
We evaluate the diffusion model guided by each reward function in a single-agent environment. The environment contains an obstacle moving from location $(10,0)$ to $(0,0)$ (right to left). The robot aims to reach location $(10,0)$ from $(0,0)$ (left to right) while avoiding the oncoming obstacle. The position and velocity of the obstacle are provided in advance for generating a plan for the robot. The initial trajectory given at the beginning of reverse diffusion is a straight line connecting the given start and goal points, and is therefore infeasible as it collides with the obstacle.
Table I summarizes the results of each method, and Fig. 3 shows the trajectory of each simulation. Note that the CBF-QP and VO methods are deterministic. CoBL generated collision-free plans that were the most smooth. In terms of goal-reaching performance, CoBL was comparable with the other CoBL variants; dCoL had the best goal-reaching performance but was worse in collision avoidance and smoothness metrics. The plans generated by CBF-QP and VO were successful in avoiding obstacles, but were significantly less smooth. VO had the worst goal-reaching performance. Ablation studies show that the safety of the generated plans deteriorates when discrete-time rewards are used to adapt to the dynamics of the system. Fig. 3 suggests that dCoBL generates more aggressive plans. This behavior is likely due to the stronger influence of the CLF in the discrete-time formulation. Therefore, it was confirmed decreasing the coefficient of discrete-time CLF reward improved safety performance. Comparing CoBL and DCoL, using a reward function derived from CBF ensures stronger safety than merely considering the distance from obstacles. When comparing CoBL with CoB, the contribution of the CLF reward to the goal convergence performance is not evident in the simple environment. Even without the CLF reward, the proposed diffusion model’s ability to synthesize a controller to achieve the given states allows it to reach the goal. We next investigate CoBL-Diffusion in a dynamic multi-agent scenario.

### v2

@@ Abstract
Equipping autonomous robots with the ability to navigate safely and efficiently around humans is a crucial step toward achieving trusted robot autonomy. However, generating robot plans while ensuring safety in dynamic multi-agent environments remains a key challenge. Building upon recent work on leveraging deep generative models for robot planning in static environments, this paper proposes CoBL-Diffusion, a novel diffusion-based safe robot planner for dynamic environments. CoBL-Diffusion uses Control Barrier and Lyapunov functions to guide the denoising process of a diffusion model, iteratively refining the robot control sequence to satisfy the safety and stability constraints. We demonstrate the effectiveness of CoBL-Diffusion using two settings: a synthetic single-agent environment and a real-world pedestrian dataset. Our results show that CoBL-Diffusion generates smooth trajectories that enable the robot to reach goal locations while maintaining a low collision rate with dynamic obstacles.
@@ I Introduction
Diffusion models [14, 15] have emerged as a powerful class of deep generative models, achieving remarkable success in generating high-quality images and synthesizing speech [16, 17], and more recently, in robot trajectory generation [18, 19, 20, 21]. However, current Diffusion-based robot trajectory generation methods still lack explicit safety assurances in dynamic environments, prohibiting their application in safety-critical scenarios involving human-robot interaction. In this work, we investigate the use of diffusion models as the foundation for robot navigation planning in dynamic multi-agent environments and present a method that integrates control-theoretic safety and stability constraints into the diffusion model framework.
@@
    We propose CoBL-Diffusion, a diffusion-based robot planner for dynamic environments that incorporates control-theoretic safety and stability constraints. The diffusion process is guided by the gradient of reward functions derived from the CBF and CLF.
@@
    Our method ensures consistency between the control sequence and the resulting states by explicitly integrating generated controls through system dynamics to obtain the states. Maintaining dynamic consistency during the diffusion steps is crucial, as the CBF and CLF reward computations in CoBL-Diffusion require the state trajectory to be consistent with the control inputs.
@@ II Related work
Diffusion Models for Planning. Diffusion models, introduced by [14] and further developed by [15], are a class of deep generative models that represent data generation as an iterative denoising process. We provide a brief overview of diffusion models in Section III-A. The Diffuser [18] represents an innovative application of diffusion models to robot trajectory generation. The Diffuser combines the planning and sampling processes into a unified framework via iterative denoising of trajectories, thus enabling the generation of plans that are adaptable to various reward structures and constraints. There are follow-up works that aim to influence the diffusion process to produce trajectories that satisfy some desired properties. Work similar in spirit to our proposed method, SafeDiffuser [32] extends the application of diffusion models to safety-critical tasks by incorporating CBFs into the denoising diffusion procedure to ensure the generated trajectories satisfy safety constraints. This CBF guarantees forward invariance of the diffusion procedure rather than the conventional forward invariance in a time domain addressed in this work. Another recent work [19] proposes guided conditional diffusion in the context of simulating human driving behaviors in driving simulators. They use signal temporal logic specifications to encode desired behaviors such as goal-reaching and staying within speed limits.
@@ III-A Diffusion Models
where $\boldsymbol{\epsilon}$ is Gaussian noise, $\boldsymbol{\epsilon}_{\theta}$ is the neural network predicting the noise, and $\bar{\alpha}_{i}=\prod_{j=1}^{i}(1-\beta_{j})$ is the cumulative product of the noise schedule. To generate new samples, one starts from noise and iteratively applies (2) for a predetermined number of iterations.
@@ III-C Control Barrier Function (CBF)
$\sup_{\mathbf{u}\in\mathcal{U}}\mathopen{}\mathclose{{}\left[L_{f}h(\mathbf{x})+L_{g}h(\mathbf{x})\mathbf{u}+\alpha(h(\mathbf{x}))}\right]\geq 0,\ \forall\mathbf{x}\in\mathcal{D},$ (9)
where $L_{X}p(x)$ is the Lie derivative of a function $p:\mathcal{P}\rightarrow\mathbb{R}$ with respect of a vector field $X$ at point $x\in\mathcal{P}$.
@@ IV-B1 U-Net Conditioning on Trajectories
To maintain dynamic consistency, our model generates only control inputs, that is, compared to (4), we have $\boldsymbol{\tau}=\mathbf{u}=[\mathbf{u}_{1},\ \ldots,\ \mathbf{u}_{T}]$, and the states are obtained by feeding the controls into the system dynamics (13) (see the blue box in Fig. 2 left), rather than simultaneously predicting control inputs and states. While Diffuser conditions the goal by replacing the terminal state, the same inpainting method cannot be applied to CoBL-Diffusion since it predicts controls only. Therefore, the states to be satisfied are fed into each resolution of the U-Net [41] (see orange boxes in Fig. 2) to condition the generated controls for the next denoising step. During the training of the proposed model, the states obtained by integrating the control inputs corrupted by noise are fed into the U-Net. This is because the states provided as conditions during the generation process are noisy, obtained by integrating noisy controls in the previous diffusion step with the system dynamics (13).
@@ IV-B3 Temporal Weighting for Covariance
The behavior of a generated trajectory depends on the sequence of control inputs by integrating the dynamics. As a result, errors in control inputs accumulate over time, meaning errors at earlier timesteps can lead to large errors in states later in the horizon. Therefore, denoising the later control sequence has little effect if the earlier control inputs have not converged. To encourage the convergence of the earlier control sequence, the covariances of the forward diffusion (1) and reverse diffusion (2) are weighted by $\mathbf{V}$ as follows:
@@
where $v_{0},\ldots,v_{T}$ are scheduled to increase monotonically with respect to the time steps.
@@ IV-C Reward Function
As introduced in Section III-B, a user-defined reward function can be used to guide the reverse diffusion process (5) to generate outputs with some desirable behavior. The reverse diffusion process can be guided by multiple reward functions simultaneously, therefore the optimality of the trajectory is denoted using the set of $K$ reward functions $W_{k}(\mathbf{x}_{t},\mathbf{u}_{t})$ as $p(O_{t}=1)=\textrm{exp}\mathopen{}\mathclose{{}\left(\sum_{k=1}^{K}W_{k}(\mathbf{x}_{t},\mathbf{u}_{t})}\right)$. Then, we modify the reverse diffusion process as follows:
@@ IV-C1 Control Barrier Function Reward
Therefore, if the reward function $W_{\mathrm{cbf}}$ is positive for each control action in the entire sequence, collision avoidance is guaranteed. The gradient of this reward function with respect to the control input $\mathbf{u}$ is computed as follows:
@@ IV-C2 Control Lyapunov Function Reward
The reward function based on Control Lyapunov Functions (CLFs) in Section III-D guides the reverse diffusion process to generate a control sequence that encourages convergence to a given goal $\mathbf{x}_{g}$. Similar to CBF reward, we assume that the dynamics of the robot is (7). Suppose $h_{\mathrm{clf}}$ is a CLF, then, according to Theorem 2, control inputs $\mathbf{u}$ that make the following reward function $W_{\mathrm{clf}}$ positive will drive the state of the robot $\mathbf{x}$ to $\mathbf{x}_{g}$:
@@
The gradient of this reward function with respect to the control input $\mathbf{u}$ is computed as follows:
@@ IV-D Conditional Motion Planning with Reverse Diffusion
First, we observe start $\mathbf{x}_{s}$ and goal $\mathbf{x}_{g}$ of the trajectory, then initialize the trajectory $\mathbf{x}^{N}$, where $N$ is the number of diffusion steps. If there is prior information about the environment, we initialize based on that; otherwise, we simply initialize the trajectory as a straight line connecting the start and goal points. Next, we sample $\mathbf{u}^{N-1}$ conditioned on $\mathbf{x}^{N}$ according to the following distribution:
@@
Then, the control inputs are fed into the system dynamics (13) to obtain the trajectory $\mathbf{x}^{N-1}$ as in line $6$ of Algorithm 1. At this point, it is important to emphasize that, unlike conventional diffusion model-based trajectory generation, the control sequence and states are consistent. Then, the trajectory is evaluated by the reward functions defined in Section IV-C. If each reward function is positive at a certain time step, it indicates the trajectory already satisfies each condition at that time. Therefore, the gradient of the reward function is masked with 00 at that point. Before fed the updated states, initial and terminal states are replaced with the given start $\mathbf{x}_{s}$ and goal $\mathbf{x}_{g}$ to condition the controls as in line $10$ of Algorithm 1. Then, the controls $\hat{\mathbf{u}}^{i}$ are denoised based on the states $\hat{\mathbf{x}}^{i}$. This approach aims to iteratively generate controls that gradually satisfy the conditions, instead of employing a common approach of solving a quadratic program (QP) to enforce satisfaction of CBF/CLF constraints which can be computationally expensive over many diffusion and timesteps. Since there is no need to solve a QP, there is no concern about the feasibility of the trajectories satisfying the constraints. When using multiple reward functions, the prioritization of conditions can be managed by adjusting their coefficients. For example, safety constraints should be emphasized over other constraints, thus necessitating a higher coefficient.
@@ V Experiments and Discussion
In this section, we validate the effectiveness of the proposed CoBL-Diffusion to synthesize safe and goal-reaching controllers in dynamic environments. First, we investigate the characteristics of different planners in a single-agent environment. Then, we validate the efficacy in a multi-agent environment using the UCY pedestrian dataset [45].
@@ V-A Model Training and Environment Setup
We trained CoBL-Diffusion using the ETH pedestrian dataset [46]; we trained for 500 epochs on 276,874 8-second trajectories. The architecture and hyperparameters of the diffusion model are based on an open-source implementation [^3] https://github.com/jannerm/diffuser. The Adam optimizer [47] was used for training with a learning rate set to 2×$\times 10^{-5}$ and $\beta=(0.9,0.999)$. The covariance weight (19) was set to increase linearly from $0.5$ to $1$. We evaluated the proposed model in two simulation environments created using Trajdata [48]. The first simulation is a toy environment with a single moving obstacle. The second environment is from the UCY pedestrian dataset [45] for challenging multi-agent yet feasible settings.
We assumed the robot follows single integrator dynamics, $\dot{\mathbf{x}}=\mathbf{u}$. We use the following CBF and CLF to guide the reverse diffusion process:
@@
where $\mathbf{x}=[x,y]$, $\mathbf{x}_{q}=[x_{q},y_{q}]$ represent the states of the robot and the $q$-th obstacle respectively, $r$ is a barrier radius, and $\mathbf{x}_{g}=[x_{g},y_{g}]$ is a given goal point.
@@ V-B Comparison Methods and Evaluation Metrics
We compare CoBL-Diffusion (CoBL) with the baseline method CBF-QP, which solves a QP with CBF at every timestep given a nominal straight line trajectory to the goal, Velocity Obstacle (VO) [49], which is a geometric approach for collision avoidance that uses the relative velocity between a robot and an obstacle to identify potential collision velocities, and Diffuser [18], which is a planning with a diffusion model and the generated plan is modified by CBF-QP to ensure collision avoidance. Variants of the proposed CoBL-Diffusion are also investigated: dCoBL-Diffusion (dCoBL), which uses discrete-time CBF and CLF rewards; CoB-Diffusion (CoB), which only uses a CBF reward; DCoL-Diffusion (DCoL), which uses CLF and a distance reward $W_{\mathrm{dist}}$ instead of a CBF reward:
@@
where $r$ is a barrier radius. Comparing the CBF reward, the distance reward considers only the distance to obstacles and ignores the dynamics. The coefficients for (discrete-time) CBF and CLF rewards are fixed to $0.3$ and $0.1$, and for distance reward is fixed to $0.3$. In this experiment, we set the barrier radius to $1\ m$ and the collision radius to $1\ m$. For a clear comparison, the reward functions for collision avoidance consider only the nearest obstacle, although considering multiple obstacles could potentially improve performance and is deferred for future work.
Each method is evaluated based on three metrics: collision rate, goal-reaching, and smoothness. The collision rate is calculated as the ratio of simulations that violate the collision radius among all simulations. The goal-reaching is evaluated by the root squared error between the goal and the final point. The smoothness is measured by the worst-case root squared difference in control inputs (in these experiments, velocity).
@@ V-C Safe Planning in Single-Agent Environment
CoBL $W_{\mathrm{CBF}}$ $W_{\mathrm{CLF}}$ $\mathbf{0}$ $\mathbf{0.20\pm 0.10}$ $\mathbf{0.45\pm 0.16}$ $\mathbf{0.5}$ $\mathbf{0.41\pm 0.20}$ $\mathbf{0.65\pm 0.31}$
CoB $W_{\mathrm{CBF}}$ $6$ $1.36\pm 0.51$ $2.01\pm 0.19$ $20.0$ $3.60\pm 1.74$ $1.66\pm 0.45$
dCoBL $W_{\mathrm{dCBF}}$ $W_{\mathrm{dCLF}}$ $8$ $0.18\pm 0.08$ $0.53\pm 0.15$ $48.0$ $0.38\pm 0.16$ $0.45\pm 0.14$
DCoL $W_{\mathrm{dist}}$ $W_{\mathrm{CLF}}$ $4$ $0.13\pm 0.06$ $0.60\pm 0.24$ $42.5$ $1.10\pm 1.89$ $1.52\pm 1.90$
CoBL- $W_{\mathrm{CBF}}$ $W_{\mathrm{CLF}}$ $8.5$ $0.48\pm 0.34$ $0.73\pm 0.39$
CBF-QP 00 % $0.13$ $10.78$ $62.5$ $0.07\pm 0.02$ $4.98\pm 4.20$
@@
(b) CBF-QP and VO.
(c) dCoBL-Diffusion.
(d) CoB-Diffusion.
(e) DCoL-Diffusion.
(f) Diffuser.
Figure 3: Generated trajectories in the single-agent environment with various planning methods. Red trajectories experience collisions.
We evaluate the diffusion model guided by each reward function in a single-agent environment. The environment contains an obstacle moving from location $(10,0)$ to $(0,0)$ (right to left). The robot aims to reach location $(10,0)$ from $(0,0)$ (left to right) while avoiding the oncoming obstacle. The position and velocity of the obstacle are provided in advance for generating a plan. The initial trajectory given at the beginning of reverse diffusion is a straight line connecting the given start and goal points, and is therefore infeasible as it collides with the obstacle.
The results of simulations are summarized in Table I, with the trajectories illustrated in Fig. 3. Note that the CBF-QP and VO methods are deterministic. The CoBL successfully generated collision-free and smooth plans. While the CBF-QP and VO methods successfully avoided obstacles, they produced less smooth trajectories, with VO also exhibiting the poor goal-reaching performance. Although the Diffuser avoided collisions with CBF-QP, it resulted in the worst deviation from the goal. Ablation studies revealed that using discrete-time rewards to adapt to system dynamics resulted in a decline in the safety of the generated plans. In terms of goal-reaching performance, CoBL was comparable with the other CoBL variants; DCoL achieved the highest goal-reaching performance but was outperformed by CoBL in terms of collision avoidance and smoothness. Fig. 3 indicates that dCoBL tends to generate more aggressive trajectories, likely due to the stronger influence of the CLF in the discrete-time formulation. Comparing CoBL and DCoL, using a reward function derived from CBF ensures stronger safety than merely considering the distance from obstacles. When comparing CoBL with CoB, the contribution of the CLF reward to the goal convergence performance is not evident in the simple environment. Even without the CLF reward, the proposed diffusion model’s ability to synthesize a controller to achieve the given states allows it to reach the goal. We next investigate CoBL-Diffusion in a dynamic multi-agent scenario.